OpenStudy (anonymous):

2700=300*2^(10x) , solve for x.

OpenStudy (anonymous):

similar to your prev question but use log instead of ln :)

OpenStudy (anonymous):

not exactly you need the change of base formula for this one

OpenStudy (anonymous):

how?

OpenStudy (campbell_st):

this is another log question divide both sides by 300 and you get \[9 = 2^{10x}\] so the question is basically asking 2 to what power is 9 so you will need to take the log of both sides... base e or base 10... doesn't matter \[\log(9) = \log(2^{10x})\] there is a log law that applies to powers \[\log(a^b) = b \times \log(a)\] apply the law and then you'll be able to solve for x

OpenStudy (anonymous):

divide both sides by 300 first b/f using log.

OpenStudy (anonymous):

\[b^x=A\iff x=\frac{\ln(A)}{\ln(b)}\] you have \[9=2^{10x}\iff 10x=\frac{\ln(9)}{\ln(2)}\]

OpenStudy (anonymous):

x=log(9) __________ 10*log(2) correct?

OpenStudy (anonymous):

i thought LN was used only with e ?

OpenStudy (anonymous):

yes that is correct, what you have change of base and of course a calculator if you want the numeric answer

OpenStudy (anonymous):

I also get x=ln(9) ________ 10*ln(2)

OpenStudy (campbell_st):

the answer will be the same... if you use base 10 or base e logs..... it will also be the same as long as the use the same base log in the numerator and denominator